# Crystal-LSBO: Automated Design of De Novo Crystals with Latent Space Bayesian Optimization

### v1

@@ 1 Introduction
The field of generative models for crystal design has emerged as a focal point for researchers from diverse disciplines. Recent literature highlights various efforts employing techniques such as diffusion models [24], generative adversarial networks (GANs) [13], autoencoders (AEs) [17], and variational autoencoders (VAEs) [20]. Each of these methods addresses the challenge of capturing the complex nature of crystals, which encompasses factors like lattice structures, types of elements, and their coordinates. This complexity contributes to an expansive search space for designing new crystals with desired properties. Consequently, a generative model capable of learning lower-dimensional representations of input instances proves invaluable. Such a model can significantly reduce the search space, enabling more efficient exploration for target crystals with specific features. Among the aforementioned methods, VAEs stand out for their ability to map complex crystal structures into a manageable, lower-dimensional latent space and generate new instances from it, enabling efficient search space for de novo crystal designs. Thus, our study employs a VAE-based approach, leveraging VAEs’ potential and their effective integration with Bayesian Optimization (BO) techniques, referred to as Latent Space Bayesian Optimization (LSBO) [7]. LSBO operates in the latent space of VAE to balance exploration and exploitation efficiently, with demonstrated success in the discovery of novel instances in diverse fields. However, the existing straightforward application of LSBO encounters difficulties in the exploration phase with the complexity of crystal structures, typically focusing only on exploitation, i.e., being able to generate valid outputs only within the immediate vicinity of latent representations of known crystals. Given the considerable time and costs involved in crystal design, there is an urgent need for a sample-efficient methodology that enables extensive exploration of de novo crystals.
@@
To showcase the effectiveness of the Crystal-LSBO framework, we applied it to design de novo crystals with optimal formation energy values, a property that is related to the overall stability and functionality of materials. Results show the effectiveness of Crystal-LSBO, paving the way for new advancements in material science. Our contributions are listed as follows:
@@
    We apply LSBO to crystal design through the Crystal-LSBO framework, validating its effectiveness through de novo design experiments focusing on formation energy values.
@@ 2 Related Works
In crystal generative modeling, the intricate nature of crystal structure representations often requires employing complex generative models or using multiple models. For instance, Hoffmann et al. [11] employed a VAE paired with a U-Net to learn from 3D density maps, and Court et al. [6] proposed a Conditional-VAE and U-Net pairing. Another approach explored by Noh et al. [17] utilizes two AEs and a VAE, where embeddings of lattice parameters and chemical elements are learned by separate AEs and used as input to the VAE model to form the generative model. Besides, the literature offers diverse methods for representing crystal structure data. Chiba et al. [5] introduced neural structure fields for use in AE-based architectures. Ren et al. [19] developed an invertible representation combining real and reciprocal space features for VAE training. Building on this, Ren et al. [20] refined this method by incorporating Fourier-transformed crystal properties (FTCP) into the VAE, referred to as FTCP-VAE. Diverging from VAEs, GANs also found themselves applications in crystal generation tasks with examples like [13, 25]. Besides, in [24], a diffusion model based crystal generation and optimization method is proposed, and Xie et al. [23] employed a diffusion model in conjunction with a VAE. Although generative models and representation techniques of the crystals differ among these studies, the common strategy for generating new crystals is similar, which is to sample from the close neighborhood of the known crystals in the latent space. This is mainly because invalid crystals are generated if we sample from a wide range across the latent space.
@@ 3.5 Latent Consistent-Aware LSBO
In order to enhance exploration in the latent space of LSBO, we adopt the Latent Consistent-Aware LSBO (LCA-LSBO) [1], recently proposed in the field of organic molecule generation. Retraining of the VAE discussed above involves updating the VAE’s training dataset with new instances from LSBO queries, and periodically retraining the VAE with this updated dataset. However, due to the expensive nature of BB function evaluations, it is unrealistic to expect a high number of new instances that are enough to bring a meaningful update to the VAE. LCA-LSBO uses label-free data augmentations in the latent space to address this challenge. Synthetic latent variables $\hat{\bm{z}}$ are generated from a probability distribution. These augmentations focus on regions of interest in the latent space that decode into instances with high property values or areas identified as promising by the AF of LSBO. Augmented variables are decoded, re-encoded, and discrepancies are penalized during VAE retraining. Namely, given augmented latent variable $\bm{\hat{z}}$, LCA-LSBO adds the additional term $||\bm{\hat{z}}-f^{\text{enc}}_{\phi}(f^{\text{dec}}_{\theta}(\bm{\hat{z}}))||^{2}$ to the objective function of VAE during retraining. This can be considered as the reconstruction of augmented latent variables at targeted regions. It provides the effective incorporation of data augmentations into the retraining process to address the problem of limited new data to use in retraining. Centered on the region of interest, these augmentations enable LSBO to more effectively explore and exploit promising regions in the latent space. The label-free nature of these augmentations and their targeted selection accelerate de novo discovery and enhance sample efficiency. For more detail on LCA-LSBO, see the appendix §A.1 and [1].
@@ 4.1 VAE Components of the Crystal-LSBO Framework
where $w$ is the weight assigned to the property predictor. [^3] Information on the model architectures of the VAEs, the PP, and training details are provided in §A.3.
@@ 5.1 Optimizing Electronegativity for Model Selection
Figure 4: Panel A showcases the outcomes of using the Crystal-Standard-LSBO method for the LSBO task focused on designing de novo crystals with enhanced electronegativity values. The 20-dimensional latent space model emerged as optimal. Panel B focuses on the optimization of predicted formation energies, comparing the performance of Crystal-Standard-LSBO, Crystal-LCA-LSBO, Random Search, FTCP-VAE Random Search, and FTCP-VAE Local Search. Crystal-LCA-LSBO significantly outperformed the others, while Crystal-Standard-LSBO also showed effective results.
@@ 5.2 Optimizing Formation Energy
This section focuses on our main goal: designing crystals with optimal formation energy values. Due to the high resource demands of exact formation energy calculations, we employed a proxy model as our BB function for estimating the formation energies of generated crystals during LSBO. We utilized an Xgboost [4] model trained on our crystal structure data and their formation energies, achieving a test Mean Squared Error (MSE) of 0.04 in predicting formation energies. Using the selected model in §5.1, we implemented the proposed Crystal-Standard-LSBO and Crystal-LCA-LSBO algorithms [^4] Details on hyperparameter selection for Crystal-LCA-LSBO are provided in the §A.4.. The GP for each of the LSBO experiments was trained on the top 100 instances with the lowest predicted formation energies. To benchmark the effectiveness of the LSBO methods, we implemented a random search strategy using both our VAEs and the FTCP-VAE model, maintaining the same search bounds as those used in the LSBO experiments. Additionally, we adopted the local search methodology proposed by the FTCP-VAE authors, which involves sampling from the vicinity of the latent representation of the crystal with the best predicted formation energy. Each method was evaluated over 1,000 iterations across 10 different seeds.
@@
Figure 4(B) displays the averaged results and standard errors across different seeds, where lower values indicate better performance. Remarkably, both the Crystal-Standard-LSBO and Crystal-LCA-LSBO methods exceeded initial benchmarks by identifying crystals with reduced predicted formation energies consistently across all seeds. In contrast, the FTCP-VAE Local Search approach only marginally surpassed the initial benchmarks in 2 of the 10 seeds, failing to deliver consistent improvements. Furthermore, while random search strategies using our model and the FTCP-VAE model failed to provide improvements, our model demonstrated comparatively better results. These results highlight the effectiveness of our Crystal-LSBO methodology in efficiently exploring the latent space to pinpoint crystals that align with our desired criteria. Among the two proposed LSBO approaches, Crystal-LCA-LSBO showcased superior capability in locating crystals with even lower predicted formation energies. Table 1 lists the three lowest predicted formation energies obtained by each method, consolidating results from all seeds. Here, the leading values achieved by Crystal-LCA-LSBO surpass those from the other methods, emphasizing its enhanced performance.
@@ 6 Conclusion
Limitations and Future Work: As discussed in §4.1, our study uses space group 1 representations of crystals for their simplicity and comprehensiveness, which comes from omitting the symmetry information of crystals. While using this representation streamlines model training, it sacrifices a structural information which can limit the generative and de novo design capabilities of the Crystal-LSBO framework. Future work could explore integrating other space group representations into our framework for richer and more targeted exploration of the crystallographic landscape.

### v2

@@ 1 Introduction
The field of generative models for crystal design has emerged as a focal point for researchers from diverse disciplines. Recent literature highlights various efforts employing techniques such as diffusion models [23, 24], generative adversarial networks (GANs) [13], autoencoders (AEs) [17], and variational autoencoders (VAEs) [20]. Each of these methods addresses the challenge of capturing the complex nature of crystals, which encompasses factors like lattice structures, types of elements, and their coordinates. This complexity contributes to an expansive search space for designing new crystals with desired properties. Consequently, a generative model capable of learning lower-dimensional representations of input instances proves invaluable. Such a model can significantly reduce the search space, enabling more efficient exploration for target crystals with specific features. Among the aforementioned methods, VAEs stand out for their ability to map complex crystal structures into a manageable, lower-dimensional latent space and generate new instances from it, enabling efficient search space for de novo crystal designs. Thus, our study employs a VAE-based approach, leveraging VAEs’ potential and their effective integration with Bayesian Optimization (BO) techniques, referred to as Latent Space Bayesian Optimization (LSBO) [7]. LSBO operates in the latent space of VAE to balance exploration and exploitation efficiently, with demonstrated success in the discovery of novel instances in diverse fields. However, the existing straightforward application of LSBO encounters difficulties in the exploration phase with the complexity of crystal structures, typically focusing only on exploitation, i.e., being able to generate valid outputs only within the immediate vicinity of latent representations of known crystals. Given the considerable time and costs involved in crystal design, there is an urgent need for a sample-efficient methodology that enables extensive exploration of de novo crystals.
@@
To showcase the effectiveness of the Crystal-LSBO framework, we applied it to design de novo crystals with optimal formation energy values, a property that is related to the overall stability and functionality of materials. Results show the effectiveness of Crystal-LSBO, paving the way for new advancements in material science. The implementation of the Crystal-LSBO framework is available at https://github.com/onurboyar/Crystal-LSBO. Our contributions are listed as follows:
@@
    We apply LSBO to crystal design through the Crystal-LSBO framework, demonstrating its effectiveness through experimental evaluations focused on formation energy. Our results are benchmarked against existing methodologies, showcasing significant improvements.
@@ 2 Related Works
In crystal generative modeling, the intricate nature of crystal structure representations often requires employing complex generative models or using multiple models. For instance, Hoffmann et al. [11] employed a VAE paired with a U-Net to learn from 3D density maps, and Court et al. [6] proposed a Conditional-VAE and U-Net pairing. Another approach explored by Noh et al. [17] utilizes two AEs and a VAE, where embeddings of lattice parameters and chemical elements are learned by separate AEs and used as input to the VAE model to form the generative model. Besides, the literature offers diverse methods for representing crystal structure data. Chiba et al. [5] introduced neural structure fields for use in AE-based architectures. Ren et al. [19] developed an invertible representation combining real and reciprocal space features for VAE training. Building on this, Ren et al. [20] refined this method by incorporating Fourier-transformed crystal properties (FTCP) into the VAE, referred to as FTCP-VAE. Although these studies employ encoding-based approaches, the high dimensionality of the latent spaces makes effective search and optimization challenging. Diverging from VAEs, GANs also found themselves applications in crystal generation tasks with examples like [13, 25]. Besides, in [24], a diffusion model based crystal generation and optimization method is proposed, and Xie et al. [23] employed a diffusion model in conjunction with a VAE. Although diffusion models demonstrate promising results in crystal design, their low sample efficiency due to high dimensional search space and the inherent stochasticity of the diffusion process hinder their effectiveness in de novo design tasks.
Despite variations in generative models and crystal representation techniques across the aforementioned studies, a common strategy for generating new crystals is to sample from the vicinity of known crystals in the latent space. This approach is primarily adopted for two reasons: first, the vast search space resulting from high-dimensional latent representations makes it challenging to explore effectively; second, sampling broadly across the latent space often leads to the generation of invalid crystal structures.
@@ 3 Preliminaries and Problem Setup
In this section, we provide preliminary knowledge on crystals, VAEs, and LSBO. We then discuss the challenges of property optimization and material space exploration using generative models.
3.1 Crystals
Crystals are characterized by an organized arrangement of elements, forming a three-dimensional lattice that repeats consistently throughout the material. This repeating structure is constructed from a basic unit called the unit cell, which acts as the fundamental building block of the crystal. The unique arrangement of elements within these unit cells dictates the material’s physical and chemical properties.
Below are definitions of key crystal-related terms used in this paper:
  • Lattice: A lattice is a 3-dimensional framework that shows where atoms are located in a crystal. The structure of the lattice is defined by a set of vectors, called lattice vectors, along with their lengths and the angles between them, which together form the entire crystal structure. The lattice serves as the framework that organizes the atomic arrangement of the material.
  • Space Groups: Space groups categorize the patterns and symmetry in crystal structures, formed by operations like translations, rotations, reflections, and inversions. There are 230 space groups, each defining unique symmetry operations and constraints about how the elements within the crystal and the lattice should be arranged [2]. Space group 1 is the simplest, involving only translations, meaning the pattern repeats periodically without additional symmetry. This group has minimal geometric constraints, allowing flexible lattice angles and lengths, and all crystals can be represented in this form by simplifying their original symmetry.
  • Primitive Cell: The primitive cell is the smallest possible unit cell that, when repeated in the crystal space, fully describes the entire crystal structure. It contains only one lattice and represents the simplest building block of the crystal. The primitive cell is unique in that it has the minimum volume and includes all the symmetry and properties of the crystal lattice.
  • Coordinates: Coordinates refer to the specific positions of atoms within the unit cell or primitive cell, typically expressed in terms of fractions of the lattice vectors. These fractional coordinates describe where each atom is located relative to the origin of the unit cell and are crucial for determining the exact structure of the crystal.
  • Elements within the Material: The elements within the material refer to the specific types of atoms that make up the crystal or material structure. These elements are defined by their atomic types (such as carbon, oxygen, hydrogen, etc.) and are fundamental to the chemical composition of the material. The arrangement and interaction of these elements within the crystal lattice determine the material’s physical and chemical properties. Figure 2(A) demonstrates the components of the crystals.
@@ 3.6 Latent Consistent-Aware LSBO
In order to enhance exploration in the latent space of LSBO, we adopt the Latent Consistent-Aware LSBO (LCA-LSBO) [1], recently proposed in the field of organic molecule generation. Retraining of the VAE discussed above involves updating the VAE’s training dataset with new instances from LSBO queries, and periodically retraining the VAE with this updated dataset. However, due to the expensive nature of BB function evaluations, it is unrealistic to expect a high number of new instances that are enough to bring a meaningful update to the VAE. LCA-LSBO uses label-free data augmentations in the latent space to address this challenge. Synthetic latent variables $\hat{\bm{z}}$ are generated from a probability distribution. These augmentations focus on regions of interest in the latent space that decode into instances with high property values or areas identified as promising by the AF of LSBO. Augmented variables are decoded, re-encoded, and discrepancies are penalized during VAE retraining. Namely, given augmented latent variable $\bm{\hat{z}}$, LCA-LSBO adds the additional term $||\bm{\hat{z}}-f^{\text{enc}}_{\phi}(f^{\text{dec}}_{\theta}(\bm{\hat{z}}))||^{2}$ to the objective function of VAE during retraining. This can be considered as the reconstruction of augmented latent variables at targeted regions. It provides the effective incorporation of data augmentations into the retraining process to address the problem of limited new data to use in retraining. Centered on the region of interest, these augmentations enable LSBO to more effectively explore and exploit promising regions in the latent space. The label-free nature of these augmentations and their targeted selection accelerate de novo discovery and enhance sample efficiency. For more detail on LCA-LSBO, see the Appendix and [1].
@@ 4.1 VAE Components of the Crystal-LSBO Framework
where $w$ is the weight assigned to the property predictor. [^3] Information on the model architectures of the VAEs, the PP, and training details are provided in the Appendix.
@@ 5.1 Optimizing Electronegativity for Model Selection
Figure 4: Panel A showcases the outcomes of using the Crystal-Standard-LSBO method for the LSBO task focused on designing de novo crystals with enhanced electronegativity values. The 20-dimensional latent space model emerged as optimal. Panel B focuses on the optimization of predicted formation energies, comparing the performance of Crystal-Standard-LSBO, Crystal-LCA-LSBO, CD-VAE, Random Search, FTCP-VAE Random Search, and FTCP-VAE Local Search. Crystal-LCA-LSBO significantly outperformed the others, while Crystal-Standard-LSBO also showed effective results.
@@ 5.2 Optimizing Formation Energy
This section focuses on our main goal: designing crystals with optimal formation energy values. Due to the high resource demands of exact formation energy calculations, we employed a proxy model as our BB function for estimating the formation energies of generated crystals during LSBO. We utilized an Xgboost [4] model trained on our crystal structure data and their formation energies, achieving a test Mean Squared Error (MSE) of 0.04 in predicting formation energies. Using the selected model in §5.1, we implemented the proposed Crystal-Standard-LSBO and Crystal-LCA-LSBO algorithms [^4] Details on hyperparameter selection for Crystal-LCA-LSBO are provided in the Appendix.. The GP for each of the LSBO experiments was trained on the top 100 instances with the lowest predicted formation energies. To benchmark the effectiveness of the LSBO methods, we implemented a random search strategy using both our VAEs and the FTCP-VAE model, maintaining the same search bounds as those used in the LSBO experiments. Additionally, we adopted the local search methodology proposed by the FTCP-VAE authors, which involves sampling from the vicinity of the latent representation of the crystal with the best-predicted formation energy. We also applied the CD-VAE methodology proposed by [23], which uses a crystal diffusion model and incorporates a VAE to condition the generation process. Their approach uses gradient-based optimization, guided by a property predictor to steer the generation towards desired properties [^5] The code provided by the CD-VAE authors was used.. Each method was evaluated over 1,000 iterations across 10 different seeds.
@@
Figure 4(B) presents the average results and standard errors across different seeds, with lower values indicating better performance. Both the Crystal-Standard-LSBO and Crystal-LCA-LSBO methods consistently outperformed initial benchmarks by identifying crystals with lower predicted formation energies across all seeds. In contrast, the FTCP-VAE Local Search approach only marginally surpassed the initial benchmarks in 2 of the 10 seeds, failing to deliver consistent improvements. Furthermore, while random search strategies using our model and the FTCP-VAE model failed to provide improvements, our model demonstrated comparatively better results. CD-VAE results, on the other hand, show clear improvements over random search approaches and slightly outperform the FTCP-VAE Local Search, which focuses on sampling near the best-known values. However, CD-VAE struggles to consistently surpass the initial benchmark values and is outperformed by the Crystal-LSBO methodologies. In each optimization experiment, CD-VAE generates 10 candidates for each of the 1000 iterations and selects the best one using an additional predictor model. This results in CD-VAE sampling 10,000 materials per experiment, compared to 1,000 samples in other models. While this extensive sampling contributes to improved performance over FTCP-VAE, it also highlights the inherent complexity and inefficiency of the search and optimization process in CD-VAE. These findings underscore the efficiency of our Crystal-LSBO methodology in exploring the latent space to identify crystals that meet the desired criteria.
Among the two LSBO approaches, Crystal-LCA-LSBO proved particularly effective, consistently finding crystals with lower predicted formation energies. Table 1 consolidates the three lowest predicted formation energies for each method across all seeds, with the best results achieved by Crystal-LCA-LSBO, further highlighting its superior performance.